OpenStudy (anonymous):

I need help on solving three system of equations, and finding the three variables. I also need to know the steps of doing it. using Elimination -3x+3y+4z=0 -3x-y+z=-3 3x+4y+4z=5

OpenStudy (anonymous):

I really need help...:(

OpenStudy (campbell_st):

the easiest thing to do is add the 1st and last equations then 2nd and last equations.... this will eliminate x and give you 2 equations in 2 unknowns. -3x + 3y + 4z = 0 Equation 1 -3x - y + z = -3 Equation 2 3x + 4y + 4z = 5 Equation 3 so Equation 1 + Equeation 3 -3x + 3y + 4z = 0 3x + 4y + 4z = 5 ------------------ 7y + 8z = 5 Equation 4 repeat the process, Equation 2 + Equation 3 -3x - y + z = -3 3x +4y + 4z = 5 --------------- 3y + 5z = 2 Equation 5 so you now have 2 equations in 2 unknowns 7y + 8z = 5 3y + 5z = 2 you will need to choose a letter to eliminate and carry out the appropriate multiplication and addions/subtractions.

OpenStudy (anonymous):

Thanks fot help

OpenStudy (anonymous):

So what are the three variables are